# Magnetar structure in nonlinear electrodynamics with mixed poloidal-toroidal fields

### v1

@@ Abstract
Magnetars have inferred polar field strengths in excess of the Schwinger limit, where nonlinear electromagnetic effects can be significant. Their internal fields may be even stronger, suggesting that Maxwellian characterisations of hydromagnetic structure may require revision. A generalised Grad-Shafranov equation, describing static and axisymmetric fluid stars with mixed poloidal-toroidal fields, is introduced and subsequently solved in a perturbative scheme to calculate quadrupolar deformations. In the Born-Infeld theory, we show that the toroidal field has a maximum strength set by the scale parameter, $b$, implying an upper limit to the stellar prolateness, $|\epsilon_{\rm max}|\sim 10^{-5}\left(b/10^{16}\text{ G}\right)^{2}$, that is independent of field specifics. Observations of magnetar phenomena that are interpreted as evidence for ellipticity, such as precession, can thus implicitly constrain post-Maxwellian parameters in a way that complements terrestrial experiments. Toroidal ceilings also have implications for dynamo theory and gravitational waves, which we revisit together with field evolution in crusts abiding by beyond-Maxwell physics.
@@
pagerange: Magnetar structure in nonlinear electrodynamics with mixed poloidal-toroidal fieldsApubyear: ?
@@ 2.3 Parameterised post-Maxwellian theory
$T_{\rm PM}(r,\theta)\propto\left[\alpha_{\rm PM}r^{4}\csc^{2}\theta I(\psi)^{2}-\frac{4}{27}\left(r^{2}-\alpha_{\rm PM}\csc^{2}\theta\nabla\psi\cdot\nabla\psi\right)^{3}\right]^{-1/2}.$ (16)
@@ 3.1 Force-free poloidal fields
and expand out equation (17) by reading off the coefficients of $\alpha_{\rm PM}^{k}$ for each $k$, a series solution can be found. Note that the Maxwell contribution takes the well-known form $f_{1,0}=B_{p}/2r$ for polar field strength $B_{p}$ (twice that of the equatorial value). Expanding out terms to successive orders, we find the solution
$f_{1}(r)=\frac{B_{p}}{2}\left(\frac{1}{r}+\frac{\alpha_{\rm PM}}{30}\frac{B_{p}^{2}}{r^{7}}+\frac{43\alpha_{\rm PM}^{2}}{3000}\frac{B_{p}^{4}}{r^{13}}\right)+\mathcal{O}(\alpha_{\rm PM}^{3},\alpha_{\rm PPM},\ldots),$ (21)
@@
which is $\ll 1$ even for $B_{p}^{2}\alpha_{\rm PM}\sim 1$. Note if (22) were of order unity this would signal the breakdown of our perturbative approach, either at the level of the radial expansion (20) or the action (15). The former can at least be quantified easily by examining the ratio $f_{1,2}(1)/f_{1,1}(1)=\frac{43}{100}\alpha_{\rm PM}B_{p}^{2}$. For BI theory, we report here the PPM solution
$\psi_{\rm BI,PPM}(r,\theta)=\frac{B_{p}}{2}\left(\frac{1}{r}+\frac{1}{60r^{7}}\frac{B_{p}^{2}}{b^{2}}+\frac{31}{84000r^{13}}\frac{B_{p}^{4}}{b^{4}}-\frac{5479}{7056000r^{19}}\frac{B_{p}^{6}}{b^{6}}\right),$ (23)
which will be useful for constructing internal fields in Sec. 3.2. For the Galactic magnetar SGR 1806–20 with $B_{p}\approx 4\times 10^{15}$ G (Olausen & Kaspi, 2014), we see from expressions (22) and (23) that the size of relative corrections are small with $\alpha_{\rm PM}\sim 5\times 10^{-33}\text{ G}^{-2}$ if $b\sim 10^{16}$ G. Strictly speaking, the PPM corrections are no stronger than the PM ones for $B_{p}\lesssim 0.7b$.
We close this section with some remarks on normalisation. Although we have and continue to use the symbol $B_{p}$ to refer to the polar field strength, it is not really so in NLED. This can be seen by substituting $r=1$ into expressions (21) or (23), revealing an increase to the polar cap strength. Although we could tune the respective coefficients between the Maxwell and NLED cases to match (for instance) the values $\psi(1,\pi/2)$, this would instead mean mismatching the asymptotics (i.e. the value of $\underset{r\to\infty}{\lim}r\psi$). Such an issue is familiar in studies of general-relativistic extensions of hydromagnetic structure (Rezzolla & Ahmedov, 2004; Pétri, 2015). We thus refer to $B_{p}$ as the polar strength in either case, to be interpreted as the polar dipole moment divided by the cube of the stellar radius, as measured by an observer at infinity.
@@ 3.2 Internal fields
Figure 2: Poloidal field lines associated with expression (27) at successive orders in the BI parameter $b$, as per plot legends (increasing in PM order from top-left to bottom-right). The colour scale depicts the poloidal field strength (redder shades indicating greater $|\bm{B}_{\rm pol}|$) relative to the equatorial value, $B_{\rm eq}$, for fixed $B_{\rm eq}=7b/20$, $d_{0}=0$, and $d_{i}=-1$ for $i=1,2,3$. The red annulus marks the location of a hypothetical crust, $0.9\leq r\leq 1$. At the outer boundary, the field matches continuously to the perturbative PPM solution (23).
@@
$\psi(r,\theta)\equiv f(r)\sin^{2}\theta=\frac{B_{p}}{2}\left(a_{2}r^{2}+a_{4}r^{4}+a_{6}r^{6}+dr^{8}\right)\sin^{2}\theta,$ (26)
where the minimum power ($r^{2}$) is chosen so that the current is everywhere finite. The extra constant $d\equiv a_{8}$ can be kept to explore a wider range of equilibria. In order to consistently match with an external field (23) (or any other), we further write each (dimensionless) constant as a series in inverse powers of $b$, such as $a_{2}=a_{20}+a_{21}B_{p}^{2}/b^{2}+a_{22}B_{p}^{4}/b^{4}+a_{23}B_{p}^{6}/b^{6}$, and equate coefficients at each order. We remark at this point that equation (25) is solved in subsequent sections with the full nonlinear Lagrangian $L_{\rm BI}$ (as we may have $B\gtrsim b$ for $r\ll 1$), even though we match to the perturbative solution at the magnetospheric boundary (where $B\lesssim 0.7b$; cf. Fig. 1). Imposing conditions (i) and (ii) described above, we find the lengthy expression
$f(r)=$ $\frac{B_{p}}{2}\Bigg{\{}\left[\frac{35}{8}-d_{0}+\frac{\left(\tfrac{143}{480}-d_{1}\right)B_{p}^{2}}{b^{2}}+\frac{\left(\tfrac{10013}{672000}-d_{2}\right)B_{p}^{4}}{b^{4}}-\frac{\left(\tfrac{126017}{2257920}+d_{3}\right)B_{p}^{6}}{b^{6}}\right]r^{2}$ (27)
$-\left[\frac{21}{4}-3d_{0}+\frac{\left(\tfrac{39}{80}-3d_{1}\right)B_{p}^{2}}{b^{2}}+\frac{\left(\tfrac{589}{22400}-3d_{2}\right)B_{p}^{4}}{b^{4}}-\frac{\left(\tfrac{5479}{53760}+3d_{3}\right)B_{p}^{6}}{b^{6}}\right]r^{4}$
$+\left[\frac{15}{8}-3d_{0}+\frac{\left(\tfrac{33}{160}-3d_{1}\right)B_{p}^{2}}{b^{2}}+\frac{\left(\tfrac{527}{44800}-3d_{2}\right)B_{p}^{4}}{b^{4}}-\frac{\left(\tfrac{126017}{2688000}+3d_{3}\right)B_{p}^{6}}{b^{6}}\right]r^{6}$
$+\left(d_{0}+\frac{d_{1}B_{p}^{2}}{b^{2}}+\frac{d_{2}B_{p}^{4}}{b^{4}}+\frac{d_{3}B_{p}^{6}}{b^{6}}\right)r^{8}\Bigg{\}},$
@@
Although the above configurations are in some sense arbitrary because they have not been derived from the true equations of motion (24), they are qualitatively similar to cases which are. This is demonstrated in Appendix A showing the ‘true’ solution for a constant-density star, where the leading-order ($b^{2}$) corrections lead to a similar increase in $B_{\rm max}$ and toroidal widening.
@@ 4 Quadrupolar ellipticity
Figure 3: Quadrupolar ellipticity (34) as a function of the magnetic energy partition $\Lambda$ (35) for the Maxwell (dashed, black) and BI (solid, red) magnetic configurations (27) for $B_{p}=10^{14}$ G. For either theory, the ellipticity vanishes at $\Lambda=\Lambda_{\rm c}\approx 0.351$, with poloidally-dominated ($\Lambda>\Lambda_{\rm c}$) and toroidally-dominated ($\Lambda<\Lambda_{\rm c}$) configurations corresponding to oblate ($\epsilon>0$) and prolate ($\epsilon<0$) stars, respectively. The minimum partition, such that the BI toroidal function (13) is real, is $\Lambda_{\rm min}\approx 2.5\times 10^{-4}$.
@@
which is a function of the parameter $i_{0}$ introduced in expression (10). The ratio $\Lambda$ is also a function of $B_{p}$ in general, as the problem is no longer scale-invariant in NLED. In practice, we define a grid of $i_{0}$ values for a given $\psi$, integrate expression (34) after finding $\delta\rho$ for each $i_{0}$, and then infer the overall ellipticity profile as a function of the energy partition $\Lambda$. Note that $\Lambda=1$ corresponds to a purely poloidal field while $\Lambda\ll 1$ is toroidally-dominated. Since the toroidal (poloidal) field generally leads to prolate (oblate) distortions, the limitation (14) implies a minimum $\Lambda$, $\Lambda_{\rm min}$, and a maximum prolateness of order ($b\gg B_{p}$)
$\epsilon_{\rm max}^{\rm prolate}\approx-6\times 10^{-6}(b/10^{16}\text{ G})^{2}\left(M_{\star}/1.4M_{\odot}\right)^{-2}\left(R_{\star}/10\text{ km}\right)^{4},$ (36)
@@ 5.1 Hall-Ohm evolution
$0=\frac{\partial\bm{B}}{\partial t}-\nabla\times\left\{\frac{c^{2}\nabla\times\left[L^{\prime}(\tfrac{1}{2}\bm{B}^{2})\bm{B}\right]}{4\pi\sigma_{0}}+\frac{c\nabla\times\left[L^{\prime}(\tfrac{1}{2}\bm{B}^{2})\bm{B}\right]\times\bm{B}}{4\pi n_{e}e}\right\}.$ (42)
One sees from equation (42) that the Hall and Ohm timescales should increase relative to their Maxwellian counterparts by a factor $\sim\sqrt{1+\bm{B}^{2}/b^{2}}$. This implies that ultra-strong fields in magnetars may be preserved for longer, though the effects are minimal unless $\bm{B}^{2}\gtrsim b^{2}$. An interesting aspect of (42) is that, even in the absence of the Hall term, multipolar components should decay in unison to some degree due to the nonlinear coupling. In the Maxwell case, the decay of Ohmic modes are decoupled meaning that one may expect the dipole field to be dominant at late times as it decays slowest.
@@
Because magnetic decay is stalled, Joule heating is reduced in kind by a factor $\sim L^{\prime}(F^{2})$ implying that highly-magnetised objects should appear colder. This may help to explain why the class of ‘ultra-long period’ objects (ULPs) appear to be both cold and magnetar-like, assuming they are isolated neutron stars (cf. de Ruiter et al., 2024, for a critical discussion). These sources were seen to pulse in the radio band, in a way that is phenomenologically similar to pulsars, and yet were invisible to follow-up X-ray searches (e.g., Hurley-Walker et al., 2022). If they did indeed house polar field-strengths of order $\gtrsim 10^{16}$ G, as would fit with spindown bounds and pair-production thresholds (as for GLEAM-X J162759.5–523504.3; Suvorov & Melatos, 2023), alleviating tensions with thermal limits — which pose a challenge for the magnetar interpretation — becomes easier.

### v2

# Magnetar structure in non-linear electrodynamics with mixed poloidal-toroidal fields

@@ Abstract
Magnetars have inferred polar field strengths in excess of the Schwinger limit, where non-linear electromagnetic effects can be significant. Their internal fields may be even stronger, suggesting that Maxwellian characterizations of hydromagnetic structure may require revision. A generalized Grad-Shafranov equation, describing static and axisymmetric fluid stars with mixed poloidal-toroidal fields, is introduced and subsequently solved in a perturbative scheme to calculate quadrupolar deformations. In the Born-Infeld theory, we show that the toroidal field has a maximum strength set by the scale parameter, $b$, implying an upper limit to the stellar prolateness, $|\epsilon_{\rm max}|\sim 10^{-5}\left(b/10^{16}\text{ G}\right)^{2}$, that is independent of field specifics. Observations of magnetar phenomena that are interpreted as evidence for ellipticity, such as precession, can thus implicitly constrain post-Maxwellian parameters in a way that complements terrestrial experiments. Toroidal ceilings also have implications for dynamo theory and gravitational waves, which we revisit together with field evolution in crusts abiding by beyond-Maxwell physics.
@@
pagerange: Magnetar structure in non-linear electrodynamics with mixed poloidal-toroidal fieldsBpubyear: ?
@@ 2.3 Parameterised post-Maxwellian theory
$T_{\rm PM}(r,\theta)\propto$$\Big{\{}\Big{[}\alpha_{\rm PM}r^{4}\csc^{2}\theta I(\psi)^{2}$ (16)
$-\frac{4}{27}\left(r^{2}-\alpha_{\rm PM}\csc^{2}\theta\nabla\psi\cdot\nabla\psi\right)^{3}\Big{]}^{1/2}$
$-r^{2}\alpha_{\rm PM}^{1/2}\csc\theta I(\psi)\Big{\}}^{-1/3}.$
@@ 3.1 Force-free poloidal fields
and expand out equation (17) by reading off the coefficients of $\alpha_{\rm PM}^{k}$ for each $k$, a series solution can be found. Note that the Maxwell contribution takes the well-known form $f_{1,0}=B_{p}/2r$ for polar field strength $B_{p}$ (twice that of the equatorial value). Expanding out terms to successive orders, we find the solution (see Appendix A for a derivation)
$f_{1}(r)=$ $\frac{B_{p}}{2}\left(\frac{1}{r}+\frac{\alpha_{\rm PM}}{30}\frac{B_{p}^{2}}{r^{7}}+\frac{43\alpha_{\rm PM}^{2}}{3000}\frac{B_{p}^{4}}{r^{13}}\right)$ (21)
$+\mathcal{O}(\alpha_{\rm PM}^{3},\alpha_{\rm PPM},\ldots),$
@@
which is $\ll 1$ even for $B_{p}^{2}\alpha_{\rm PM}\sim 1$. Note if (22) were of order unity this would signal the breakdown of our perturbative approach, either at the level of the radial expansion (20) or the action (15). The former can at least be quantified easily by examining the ratio $f_{1,2}(1)/f_{1,1}(1)=\frac{43}{100}\alpha_{\rm PM}B_{p}^{2}$. For BI theory, we report here the PPM solution (Appendix A)
$\psi_{\rm BI,PPM}(r,\theta)=$ $\frac{B_{p}}{2}\Big{(}\frac{1}{r}+\frac{1}{60r^{7}}\frac{B_{p}^{2}}{b^{2}}$ (23)
$+\frac{31}{84000r^{13}}\frac{B_{p}^{4}}{b^{4}}-\frac{5479}{7056000r^{19}}\frac{B_{p}^{6}}{b^{6}}\Big{)},$
which will be useful for constructing internal fields in Sec. 3.2. For the Galactic magnetar SGR 1806–20 with $B_{p}\approx 4\times 10^{15}$ G (Olausen & Kaspi, 2014), we see from expressions (22) and (23) that the size of relative corrections are small with $\alpha_{\rm PM}\sim 5\times 10^{-33}\text{ G}^{-2}$ if $b\sim 10^{16}$ G. Strictly speaking, the PPM corrections are no stronger than the PM ones for $B_{p}\lesssim 0.7b$. While a fully-numerical, grid-based method to solve (17) would hold the advantage that no such projections are necessary, the perturbative approach used here is reasonable because all Galactic magnetars satisfy this condition (though cf. Sec. 5.3). For the interior, hydromagnetic problem – discussed next – the local field may exceed $b$ and this may no longer be the case. On the other hand, since the toroidal maximum (14) is a non-perturbative result and the field structures we find using either $\mathcal{O}(b^{-4})$ or $\mathcal{O}(b^{-6})$ expansions are quantitatively similar (Sec. 3.2), we expect that any mismatches with ‘true’ solutions are subleading with respect to other approximations made here (e.g., Newtonian gravity).
We close this section with some remarks on normalisation. Although we have and continue to use the symbol $B_{p}$ to refer to the polar field strength, it is not really so in NLED. This can be seen by substituting $r=1$ into expressions (21) or (23), revealing an increase to the polar cap strength. Although we could tune the respective coefficients between the Maxwell and NLED cases to match (for instance) the values $\psi(1,\pi/2)$, this would instead mean mismatching the asymptotics (i.e. the value of $\underset{r\to\infty}{\lim}r\psi$). Such an issue is familiar in studies of general-relativistic extensions of Maxwell’s equations (Rezzolla & Ahmedov, 2004; Pétri, 2015). We thus refer to $B_{p}$ as the polar strength in either case, to be interpreted as the polar dipole moment divided by the cube of the stellar radius, as measured by an observer at infinity.
@@ 3.2 Internal fields
Figure 2: Poloidal field lines associated with expression (3.2) at successive orders in the BI parameter $b$, as per plot legends (increasing in PM order from top-left to bottom-right). The colour scale depicts the poloidal field strength (redder shades indicating greater $|\bm{B}_{\rm pol}|$) relative to the equatorial value, $B_{\rm eq}$, for fixed $B_{\rm eq}=7b/20$, $d_{0}=0$, and $d_{i}=-1$ for $i=1,2,3$. The red annulus marks the location of a hypothetical crust, $0.9\leq r\leq 1$. At the outer boundary, the field matches continuously to the perturbative PPM solution (23).
@@
$\psi(r,\theta)$ $\equiv f(r)\sin^{2}\theta$ (26)
$=\frac{B_{p}}{2}\left(a_{2}r^{2}+a_{4}r^{4}+a_{6}r^{6}+dr^{8}\right)\sin^{2}\theta,$
where the minimum power ($r^{2}$) is chosen so that the current is everywhere finite. The extra constant $d\equiv a_{8}$ can be kept to explore a wider range of equilibria. In order to consistently match with an external field (23) (or any other), we further write each (dimensionless) constant as a series in inverse powers of $b$, such as $a_{2}=a_{20}+a_{21}B_{p}^{2}/b^{2}+a_{22}B_{p}^{4}/b^{4}+a_{23}B_{p}^{6}/b^{6}$, and equate coefficients at each order. We remark at this point that equation (25) is solved in subsequent sections with the full nonlinear Lagrangian $L_{\rm BI}$ (as we may have $B\gtrsim b$ for $r\ll 1$), even though we match to the perturbative solution at the magnetospheric boundary (where $B\lesssim 0.7b$; cf. Fig. 1). Imposing conditions (i) and (ii) described above, we find the lengthy expression (again we refer to Appendix A) {widetext}
$f(r)=$ $\frac{B_{p}}{2}\Bigg{\{}\left[\frac{35}{8}-d_{0}+\frac{\left(\tfrac{143}{480}-d_{1}\right)B_{p}^{2}}{b^{2}}+\frac{\left(\tfrac{10013}{672000}-d_{2}\right)B_{p}^{4}}{b^{4}}-\frac{\left(\tfrac{126017}{2257920}+d_{3}\right)B_{p}^{6}}{b^{6}}\right]r^{2}$
$-\left[\frac{21}{4}-3d_{0}+\frac{\left(\tfrac{39}{80}-3d_{1}\right)B_{p}^{2}}{b^{2}}+\frac{\left(\tfrac{589}{22400}-3d_{2}\right)B_{p}^{4}}{b^{4}}-\frac{\left(\tfrac{5479}{53760}+3d_{3}\right)B_{p}^{6}}{b^{6}}\right]r^{4}$ (27)
$+\left[\frac{15}{8}-3d_{0}+\frac{\left(\tfrac{33}{160}-3d_{1}\right)B_{p}^{2}}{b^{2}}+\frac{\left(\tfrac{527}{44800}-3d_{2}\right)B_{p}^{4}}{b^{4}}-\frac{\left(\tfrac{126017}{2688000}+3d_{3}\right)B_{p}^{6}}{b^{6}}\right]r^{6}+\left(d_{0}+\frac{d_{1}B_{p}^{2}}{b^{2}}+\frac{d_{2}B_{p}^{4}}{b^{4}}+\frac{d_{3}B_{p}^{6}}{b^{6}}\right)r^{8}\Bigg{\}},$
@@
Although the above configurations are in some sense arbitrary because they have not been derived from the true equations of motion (24), they are qualitatively similar to cases which are. This is demonstrated in Appendix B showing the ‘true’ solution for a constant-density star, where the leading-order ($b^{2}$) corrections lead to a similar increase in $B_{\rm max}$ and toroidal widening.
@@ 4 Quadrupolar ellipticity
Figure 3: Quadrupolar ellipticity (34) as a function of the magnetic energy partition $\Lambda$ (35) for the Maxwell (dashed, black) and BI (solid, red) magnetic configurations (3.2) for $B_{p}=10^{14}$ G. For either theory, the ellipticity vanishes at $\Lambda=\Lambda_{\rm c}\approx 0.351$, with poloidally-dominated ($\Lambda>\Lambda_{\rm c}$) and toroidally-dominated ($\Lambda<\Lambda_{\rm c}$) configurations corresponding to oblate ($\epsilon>0$) and prolate ($\epsilon<0$) stars, respectively. The minimum partition, such that the BI toroidal function (13) is real, is $\Lambda_{\rm min}\approx 2.5\times 10^{-4}$.
@@
which is a function of the parameter $i_{0}$ introduced in expression (10). The ratio $\Lambda$ is also a function of $B_{p}$ in general, as the problem is no longer scale-invariant in NLED. In practice, we define a grid of $i_{0}$ values for a given $\psi$, integrate expression (34) after finding $\delta\rho$ for each $i_{0}$, and then infer the overall ellipticity profile as a function of the energy partition $\Lambda$. Note that $\Lambda=1$ corresponds to a purely poloidal field while $\Lambda\ll 1$ is toroidally-dominated. Since the toroidal (poloidal) field generally leads to prolate (oblate [^3] Although the models constructed here are static, rotation would also induce some degree of geometric oblateness. While such a contribution is necessarily aligned with the symmetry axis and thus does not result in GW emission (which forms our primary interest here), some care is needed in applying this terminology in more general cases (see, e.g., Frieben & Rezzolla, 2012).) distortions, the limitation (14) implies a minimum $\Lambda$, $\Lambda_{\rm min}$, and a maximum prolateness of order ($b\gg B_{p}$)
$\epsilon_{\rm max}^{\rm prolate}\approx-6\times 10^{-6}\left(\frac{b}{10^{16}\text{ G}}\right)^{2}\left(\frac{1.4M_{\odot}}{M_{\star}}\right)^{2}\left(\frac{R_{\star}}{10\text{ km}}\right)^{4},$ (36)
@@ 5.1 Hall-Ohm evolution
$\frac{\partial\bm{B}}{\partial t}=\nabla\times\left\{\frac{c^{2}\nabla\times\left[L^{\prime}(\tfrac{1}{2}\bm{B}^{2})\bm{B}\right]}{4\pi\sigma_{0}}+\frac{c\nabla\times\left[L^{\prime}(\tfrac{1}{2}\bm{B}^{2})\bm{B}\right]\times\bm{B}}{4\pi n_{e}e}\right\}.$ (42)
One sees from equation (42) that the Ohm and Hall timescales should increase relative to their Maxwellian counterparts by a factor $\sim\sqrt{1+\bm{B}^{2}/b^{2}}$. This implies that ultra-strong fields in magnetars may be preserved for longer, though the effects are minimal unless $\bm{B}^{2}\gtrsim b^{2}$. An interesting aspect of (42) is that, even in the absence of the Hall term, multipolar components should decay in unison to some degree due to the nonlinear coupling. In the Maxwell case, the decay of Ohmic modes are decoupled meaning that one may expect the dipole field to be dominant at late times as it decays slowest.
Tackling equation (42) in a fully-numerical setup is challenging owing to the fact that even the most stable magnetic-evolution codes are reliable only at local strengths $\lesssim 10^{16}$ G, as the Hall prefactor grows large (see, e.g., Pons & Viganò, 2019), where NLED effects begin to dominate. Moreover, without an analytic form describing a potential field, an elliptic subroutine must also be devised to sequentially feed boundary data to ensure consistent matchings at $r=1$. This is reminiscent of evolution schema involving couplings between the crust and magnetosphere where currents thread the surface, as described (for instance) by Urbán et al. (2023); techniques presented there could be useful in this direction (especially via neural networks for rapid matching; see Urbán, Stefanou, & Pons, 2025). While a special-purpose code could be built using various methods for handling stiffness (e.g. Radau), this lies beyond the scope of the present article.
@@
Because magnetic decay is stalled, Joule heating is reduced in kind by a factor $\sim L^{\prime}(F^{2})$ implying that highly-magnetised objects should appear colder. This may help to explain why (some members of) the class of ‘ultra-long period’ objects (ULPs) appear to be both cold and magnetar-like, assuming they are isolated neutron stars (cf. de Ruiter et al., 2025, for a critical discussion). These sources were seen to pulse in the radio band in a way that is phenomenologically similar to pulsars and yet, with the exception of ASKAP J1832–0911 (Wang et al., 2024), were invisible to follow-up X-ray searches (e.g., Hurley-Walker et al., 2022). If they did indeed house polar field-strengths of order $\gtrsim 10^{16}$ G, as would fit with spindown bounds and pair-production thresholds (Suvorov & Melatos, 2023), alleviating tensions with thermal limits becomes easier. Alternative pathways for ULP production that do not require NLED-level fields have been explored though (see Cooper & Wadiasingh 2024; Suvorov, Dehman & Pons 2025.